A square and a rectangle have the same are. if the perimeter of the square is 24 inches, and the width of the rectangle is 3 inches, what is the perimeter of the rectangle?
Answer by robertb(5830)

If the square has perimeter 24 inches then its edge is 6 inches. Its area is then 36 square inches. Now the width of the rectangle is 3 inches, and its length is unknown L. Its area is then 3L square inches. Since the square and the rectangle have the same area, 3L = 36, or L = 12. Therefore the perimeter of the rectangle is 2(3 + 12) = 2*15 = 30 inches.